# SymCERE: Symmetric Contrastive Learning for Robust Review-Enhanced Recommendation

### v2

@@ Abstract.
Modern recommendation systems can achieve high performance by fusing user behavior graphs (via GNNs) and review texts (via LLMs). However, this fusion faces three significant issues: (1) False Negatives in contrastive learning can degrade the training signal by penalizing similar items; (2) Popularity Bias, often encoded as embedding magnitude, can distort similarity scores; and (3) Signal Ambiguity, which arises from the conflation of objective facts with subjective sentiment in reviews. These interconnected issues can prevent models from learning users’ true preferences.
In this paper, we propose SymCERE (Symmetric SINCERE), a contrastive learning method that addresses these three issues simultaneously through its structural design. First, we introduce a symmetric application of the SINCERE loss for cross-modal alignment, which is designed to eliminate false negatives in recommendation. Second, by integrating this with L2 normalisation under a ”magnitude-as-noise” hypothesis, we aim to mitigate popularity bias by forcing the model to encode preferences primarily in the vector’s direction.
Experiments on 15 datasets from three distinct platforms (e-commerce, local reviews, and travel) demonstrate that SymCERE outperforms several strong baselines, achieving a relative improvement of up to 43.6% on NDCG@10. Furthermore, a detailed LIME analysis shows that the model learns to anchor alignment on objective, informative vocabulary (e.g., ”OEM,” ”compatible,” ”gasket”), while placing less emphasis on generic sentiment (e.g., ”good,” ”great”). This suggests that effective semantic alignment stems from understanding factual product attributes, offering a path toward more accurate recommendation systems. The code is available at:
https://anonymous.4open.science/r/ReviewGNN-2E1E.
Recommender Systems, Contrastive Learning, False Negatives, Popularity Bias, GNN, LLM
copyright: acmlicensedjournalyear: 2018ccs: Information systems Recommender systemsccs: Computing methodologies Machine learning
@@ 1. Introduction
Recent advances in recommender systems focus on multi-modal approaches that fuse user behavior, typically captured by Graph Neural Networks (GNNs), with side information from modalities like review texts, processed by Large Language Models (LLMs) (Zhao et al., 2023; Ren et al., 2024). While this fusion can yield richer user and item representations (Cheng et al., 2024; Luo et al., 2024), its practical implementation faces a triad of interconnected challenges that can impair recommendation quality. These are: the false negative problem in contrastive learning, item popularity bias, and signal ambiguity in review texts. These issues often compound one another: popularity bias can exacerbate the false negative problem, while both can be masked by ambiguous textual signals, making it difficult to learn a user’s true intent.
While various methods have been proposed to tackle these issues individually—using denoising modules (Lin et al., 2021, 2023a) or causal inference (Zhang and Chen, 2021; Zheng et al., 2021)—we posit that these challenges can be systematically mitigated by modifying the core contrastive learning objective itself. We propose SymCERE (Symmetric SINCERE), a unified methodology that addresses this triad with geometric mechanisms. SymCERE employs a symmetric, denoised contrastive loss based on SINCERE (Zhong et al., 2023; Feeney et al., 2023) to eliminate intra-class repulsion and false negatives. Guided by our ”magnitude-as-noise” hypothesis, it integrates this with L2 normalisation to mitigate popularity bias by forcing the model to encode preferences solely in the vector’s direction.
Our experiments on 15 datasets from three distinct platforms, multiple domains, and two languages show that SymCERE significantly outperforms strong baselines, achieving up to a 43.6% relative improvement in NDCG@10. Furthermore, our analysis reveals that the model learns to anchor its alignment on objective, factual vocabulary (e.g., ‘OEM’, ‘compatible’) over generic sentiment. This mechanism, which we term ‘semantic anchoring’, suggests that robust alignment stems from understanding factual product attributes, presenting a path toward more accurate and interpretable recommender systems. Our main contributions are:
@@
  • A Unified Contrastive Method (SymCERE): We propose a method that structurally mitigates false negatives, popularity bias, and signal ambiguity by symmetrically applying a denoised SINCERE loss and integrating geometric debiasing via L2 normalisation.
@@
  • Extensive Empirical Validation: We demonstrate the effectiveness of SymCERE through comprehensive experiments on 15 diverse datasets, achieving significant improvements over strong baselines.
@@
  • A Novel Interpretation of Multi-Modal Alignment: We find, through in-depth LIME analysis, that our model learns to prioritize product-specific terminology over generic sentiment, a mechanism we term ‘semantic anchoring’.
@@ 2. Related Work
Our work lies at the intersection of multi-modal recommendation, Graph Neural Networks (GNNs), and contrastive learning (CL). While GNNs like LightGCN (He et al., 2020) form the backbone of many modern recommenders, enhancing them with review texts via CL surfaces a triad of interconnected challenges.
@@ 2.1. A Triad of Interconnected Challenges
The False Negative Problem
The standard InfoNCE loss has a key limitation: by treating all non-positive samples as negatives, it may penalize ”false negatives”—semantically similar items that are not the anchor’s direct positive pair (Huynh et al., 2022; Chen et al., 2022). This ”intra-class repulsion” can negatively affect the embedding space by pushing similar items apart and hindering the development of a coherent semantic structure. This is especially problematic in recommendation, where a user who bought one camera might also be interested in a compatible lens—an item that standard InfoNCE would treat as a hard negative.
Popularity Bias as a Geometric Artifact
Popularity bias, the over-recommendation of popular items, is a known challenge to fairness and diversity (Klimashevskaia et al., 2024). We approach this from a geometric perspective, positing that popularity is often encoded as the magnitude (L2-norm) of an item’s embedding vector (Lin et al., 2022, 2023b). This magnitude can act as a noisy, confounding signal that obscures user preferences for specific attributes. For instance, a model may incorrectly learn to equate a large embedding magnitude with high item quality, when it merely reflects high interaction frequency. Therefore, we treat L2 normalisation not just as a technical step, but as a debiasing mechanism that can structurally remove the popularity signal, building on the theoretical properties of CL on a unit hypersphere (Wang and Isola, 2020).
Signal Ambiguity in Multi-Modal Fusion
Fusing collaborative signals with review texts can enrich representations but may introduce ”signal ambiguity.” A review can conflate objective product attributes (e.g., ”OEM part,” ”compatible gasket”) with subjective, often uninformative sentiment (e.g., ”good,” ”great”) (Wu et al., 2014). A simple alignment may lead a model to learn spurious correlations from generic sentiment rather than from true semantic understanding, as it cannot distinguish the nuanced reasons for a user’s choice.
@@ 2.2. Our Approach: A Unified Solution via Principled CL
The literature often presents methods that address these issues individually (Lou et al., 2024; Lin et al., 2023a, 2024). We diverge from this trend, arguing that modifying the core CL objective itself offers a more direct and unified solution. To this end, we build upon SINCERE (Zhong et al., 2023; Feeney et al., 2023), a loss that eliminates the intra-class repulsion found in other contrastive methods. By integrating a symmetric SINCERE loss with L2 normalisation, SymCERE is designed to simultaneously: (1) Eliminate False Negatives by definition, (2) Mitigate Popularity Bias structurally, and (3) Resolve Signal Ambiguity by focusing alignment on stable, objective information. By embedding these solutions directly into the learning objective, our approach avoids the need for complex, multi-stage pipelines or heuristic post-processing. We posit that this unified approach leads to a simpler, yet more powerful and interpretable recommender system.
@@ 3. Methodology
Our proposed model architecture is designed to create a unified representation space where embeddings from user-item interactions (graph modality) and review texts (text modality) are meaningfully aligned. This is achieved through a multi-component system, trained end-to-end, which consists of modality-specific encoders and a multi-task contrastive learning objective.
@@ 3.1. Modality Encoders
We employ two specialized encoders to process the distinct data types, capturing both collaborative filtering patterns and deep semantic information.
3.1.1. Graph Encoder for Collaborative Signal
To capture the collaborative filtering signals from the user-item interaction graph, we adopt the architecture of LightGCN (He et al., 2020) as our graph encoder, denoted $f_{\text{graph}}(\cdot)$. This choice is motivated by its effectiveness in distilling the core mechanism of GNNs for recommendation—neighborhood aggregation—while removing components like feature transformations and non-linearities, which have been shown to be less critical for this task (He et al., 2020).
The encoder begins with an initial embedding layer, which provides zero-th layer embeddings $\mathbf{e}_{u}^{(0)}$ for each user $u$ and $\mathbf{e}_{i}^{(0)}$ for each item $i$. These are the only trainable parameters in the graph encoder. The embeddings are then refined over $K$ layers using a linear propagation rule that aggregates information from neighbors (Kipf and Welling, 2017):
(1) $\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)},\quad\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{i}||\mathcal{N}_{u}|}}\mathbf{e}_{u}^{(k)}$
where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with, and $\mathcal{N}_{i}$ is the set of users who have interacted with item $i$. The term $1/\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}$ is a symmetrically normalized coefficient that stabilizes the propagation process (Kipf and Welling, 2017).
The final representation for a user or item is a weighted combination of the embeddings from all layers, which captures collaborative signals from different neighborhood depths:
(2) $\mathbf{g}_{u}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{u}^{(k)},\quad\mathbf{g}_{i}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{i}^{(k)}$
where $\alpha_{k}$ are hyperparameters weighting the importance of the $k$-th layer. Following common practice, we set $\alpha_{k}=1/(K+1)$ to give equal weight to each layer (He et al., 2020). To obtain a single representation for a user-item interaction $(u,i)$, we average their final embeddings: $\mathbf{g}_{u,i}=(\mathbf{g}_{u}+\mathbf{g}_{i})/2$.
3.1.2. Text Encoder for Semantic Representation
To extract semantic information from user reviews, we employ a pre-trained Large Language Model (LLM) as our text encoder, $f_{\text{text}}(\cdot)$. We utilize Parameter-Efficient Fine-Tuning (PEFT) with LoRA (Hu et al., 2021) to adapt the LLM to the recommendation domain without the high computational cost of full fine-tuning. Given the raw review text $t_{i}$ associated with an item $i$, the encoder produces a dense semantic embedding, $\mathbf{t}_{i}=f_{\text{text}}(t_{i})$.
3.2. Unified Contrastive Learning Method
Our methodology uses a multi-task contrastive method designed to fuse the collaborative and semantic signals. It is built upon two principles: geometric debiasing and a denoised alignment objective.
3.2.1. Geometric Debiasing via L2 Normalisation
A key step, applied before any loss calculation, is the L2 normalisation of all embeddings. This is central to our ”magnitude-as-noise” hypothesis, which posits that the L2-norm of an embedding often encodes popularity bias rather than user preference (Lin et al., 2022, 2023b). By projecting all embeddings onto a unit hypersphere, we aim to eliminate this magnitude component:
(3) $\hat{\mathbf{g}}_{u,i}=\frac{\mathbf{g}_{u,i}}{\|\mathbf{g}_{u,i}\|_{2}},\quad\hat{\mathbf{t}}_{i}=\frac{\mathbf{t}_{i}}{\|\mathbf{t}_{i}\|_{2}}$
This ensures that similarity is based on angular distance (cosine similarity). This is a theoretically-grounded debiasing mechanism that satisfies the geometric precondition for contrastive learning to optimize for the properties of Alignment and Uniformity (Wang and Isola, 2020).
3.2.2. Multi-Task Contrastive Objectives
Our method is trained with a combination of a cross-modal loss for alignment and an intra-modal loss for representation enhancement (Wu et al., 2021).
@@ 3.2.2. Multi-Task Contrastive Objectives
Cross-Modal Alignment with Symmetric SINCERE Loss ($\mathcal{L}_{\text{cross-modal}}$)
To align the graph and text modalities, we address the ”false negative” problem in standard contrastive losses. The InfoNCE loss (van den Oord et al., 2018) is defined for an anchor $\mathbf{z}_{i}$, a positive sample $\mathbf{z}_{i}^{+}$, and a set of N1N-1 negative samples $\{\mathbf{z}_{j}^{-}\}_{j=1}^{N-1}$ as:
(4) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)}{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau)+\sum_{j=1}^{N-1}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{-})/\tau)}$
where $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product) and $\tau$ is a temperature parameter. This formulation treats all samples that are not the designated positive as negatives, which can lead to false negatives.
The Supervised Contrastive (SupCon) loss (Khosla et al., 2020) attempts to address this by leveraging class labels. For an anchor $\mathbf{z}_{i}$ from class $c$, it treats all other samples from class $c$ as positives. However, its formulation can lead to intra-class repulsion (Zhong et al., 2023; Feeney et al., 2023). The SupCon loss for an anchor $\mathbf{z}_{i}$ and a positive $\mathbf{z}_{p}$ from the same class is:
(5) $\mathcal{L}_{\text{SupCon}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\sum_{k\in P_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{k}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
where $P_{i}$ is the set of all positives for anchor $i$ and $N_{i}$ is the set of all negatives. The denominator includes other positive samples from the same class, which can create a repulsive gradient term (Zhong et al., 2023).
To overcome this, we adopt the SINCERE loss (Zhong et al., 2023; Feeney et al., 2023), which adheres to noise-contrastive estimation principles: only true negatives should be in the denominator (Gutmann and Hyvärinen, 2010). For an anchor $\mathbf{z}_{i}$, a positive $\mathbf{z}_{p}$, and a set of true negatives $N_{i}$, the SINCERE loss is:
(6) $\mathcal{L}_{\text{SINCERE}}=-\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)}{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{p}/\tau)+\sum_{j\in N_{i}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$
This formulation is designed to eliminate intra-class repulsion (Zhong et al., 2023). We apply this in a symmetric, cross-modal fashion. In our setting, each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph). The graph embedding $\hat{\mathbf{g}}_{u,i}$ is computed from both the user $u$ and item $i$ via the graph encoder, while the text embedding $\hat{\mathbf{t}}_{i}$ is derived from the review(s) of item $i$.
In our method, we process the data in mini-batches, where each sample in a mini-batch corresponds to a single user–item interaction (i.e., an edge in the user–item graph). For a batch of $B$ interactions, the graph-to-text loss aligns each graph embedding $\hat{\mathbf{g}}_{u,i}$ with its corresponding text embedding $\hat{\mathbf{t}}_{i}$, treating all other text embeddings $\{\hat{\mathbf{t}}_{j}\}_{j\neq i}$ as true negatives (i.e., items in the batch that have different item IDs from the anchor, excluding those the user has previously interacted with):
(7) $\mathcal{L}_{G\to T}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)}{\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{g}}_{u,i}^{\top}\hat{\mathbf{t}}_{j}/\tau)}$
Symmetrically, the text-to-graph loss is:
(8) $\mathcal{L}_{T\to G}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)}{\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,i}/\tau)+\sum_{j\in\mathcal{N}_{i}}\exp(\hat{\mathbf{t}}_{i}^{\top}\hat{\mathbf{g}}_{u,j}/\tau)}$
The final cross-modal loss is the average of these two, ensuring balanced alignment:
(9) $\mathcal{L}_{\text{cross-modal}}=\frac{1}{2}\left(\mathcal{L}_{G\to T}+\mathcal{L}_{T\to G}\right)$
Intra-Modal Alignment with InfoNCE Loss ($\mathcal{L}_{\text{intra-modal}}$)
To improve the robustness and discriminative power of the representations within each modality, we incorporate a self-supervised, intra-modal contrastive task (Wu et al., 2021). For each modality, we create two distinct ”views” through data augmentation and enforce consistency between their representations.
For the graph modality, we create an augmented view by applying random edge dropout. Let $\hat{\mathbf{g}}$ be the batch of embeddings from the original graph and $\hat{\mathbf{g}}^{\prime}$ be from the augmented graph. For the text modality, we create an augmented view via random word masking. Let $\hat{\mathbf{t}}$ be the embeddings of the original texts and $\hat{\mathbf{t}}^{\prime}$ be those of the augmented texts.
We use the standard InfoNCE loss for this task, as within-modality augmentation does not have the same false negative issue as cross-modal alignment. The loss for a generic modality with original embeddings $\hat{\mathbf{z}}$ and augmented embeddings $\hat{\mathbf{z}}^{\prime}$ is:
(10) $\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{z}},\hat{\mathbf{z}}^{\prime})=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\hat{\mathbf{z}}_{i}^{\top}\hat{\mathbf{z}}^{\prime}_{j}/\tau)}$
The total intra-modal loss is the sum of the losses from both modalities:
(11) $\mathcal{L}_{\text{intra-modal}}=\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{g}},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{t}},\hat{\mathbf{t}}^{\prime})$
This formulation can resolve signal ambiguity because by removing potential noise from false negatives (via SINCERE) and popularity signals (via L2 norm), the optimization is encouraged to find stable, information-rich correlations. In many domains, these signals are objective, factual attributes (e.g., ”OEM,” ”compatible”) rather than subjective sentiments (e.g., ”good,” ”love”), leading to a more grounded semantic alignment.
@@
In addition to aligning modalities, we regularize the embedding space within the graph modality using the BPR loss. This complements our main objective by enforcing a well-ordered ranking structure based on collaborative signals. To maintain logical consistency with our ”magnitude-as-noise” hypothesis, the BPR loss is critically applied to the L2-normalized embeddings ($\hat{\mathbf{g}}$):
(12) $\mathcal{L}_{\text{BPR}}=\sum_{(u,i,j)}-\log\sigma(\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{i}-\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{j})$

### v3

@@ Abstract.
Modern recommendation systems fuse user behavior graphs and review texts but often encounter a ”Fusion Gap” caused by False Negatives, Popularity Bias, and Signal Ambiguity. We propose SymCERE (Symmetric NCE), a contrastive learning framework bridging this gap via structural geometric alignment. First, we introduce a symmetric NCE loss that leverages full interaction history to exclude false negatives. Second, we integrate L2 normalization to structurally neutralize popularity bias. Experiments on 15 datasets (e-commerce, local reviews, travel) demonstrate that SymCERE outperforms strong baselines, improving NDCG@10 by up to 43.6%. Notably, we validate this on raw reviews, addressing significant noise. Analysis reveals ”Semantic Anchoring,” where the model aligns on objective vocabulary (e.g., ”OEM,” ”gasket”) rather than generic sentiment. This indicates effective alignment stems from extracting factual attributes, offering a path toward robust, interpretable systems. The code is available at https://anonymous.4open.science/r/ReviewGNN-2E1E.
@@ 1. Introduction
Recent advances in recommender systems focus on multi-modal approaches that fuse user behavior, typically captured by Graph Neural Networks (GNNs), with side information from modalities like review texts, processed by Large Language Models (LLMs). While this fusion promises richer user and item representations (Cheng et al., 2024a; Luo et al., 2024), its practical implementation faces a triad of interconnected challenges—a phenomenon we identify as the ”Fusion Gap”. This gap represents the disconnect between the high-quality semantic potential of LLMs and the noisy geometric reality of collaborative filtering spaces.
The first challenge contributing to this gap is the false negative problem, a known limitation in standard contrastive learning (CL) objectives like InfoNCE. By treating all non-positive pairs as negatives, standard methods inherently penalize ”false negatives”—valid positive items (whether known interactions or latent preferences) erroneously treated as negatives. This creates ”intra-class repulsion” (Chen et al., 2022; Huynh et al., 2022), pushing semantically related items apart in the embedding space and degrading the coherent structure learned by the LLM.
Figure 1. Visualizing the Fusion Gap and its Resolution. Left: Before normalization, the embedding space is dominated by magnitude artifacts (popularity bias), causing a geometric and semantic disconnect between Graph (blue) and Text (orange) modalities. Right: SymCERE structurally aligns these spaces on a unit hypersphere, enabling coherent clustering based on semantic preference rather than interaction frequency.
The second challenge is popularity bias, which acts as a geometric artifact. Popular items, due to their high interaction frequency, tend to develop embeddings with a disproportionately large L2-norm (magnitude) (Klimashevskaia et al., 2024; Jannach et al., 2021). We hypothesize that this magnitude acts as a confounding noise factor, reflecting popularity rather than a user’s intrinsic preference. When fusing modalities, this magnitude-driven noise can overshadow the delicate directional signals from the text modality (as visualized in Figure 1, Left).
The third challenge is signal ambiguity, which complicates search-to-recommendation bridging. Unlike clean product metadata (used in many MMRec tasks), a user review contains valuable semantic signals about product attributes (e.g., ”This OEM part fits perfectly.”) entangled with subjective, noisy sentiment (e.g., ”I love it!”) (Wu et al., 2014; Bauman et al., 2017). A model that cannot geometrically distinguish these signals may learn spurious correlations from generic sentiment instead of factual alignment (Cheng et al., 2024b).
While various methods have been proposed to tackle these issues individually—using denoising modules (Lin et al., 2023), causal inference for debiasing (Zhang and Chen, 2021; Zheng et al., 2021), or heuristic false-negative correction (Chen et al., 2022)—we explore a unified geometric approach. We posit that these three challenges can be systematically mitigated by modifying the core contrastive learning objective to respect the geometry of the embedding space.
We propose SymCERE (Symmetric NCE), a unified contrastive learning methodology that addresses this triad to bridge the Fusion Gap. First, to mitigate the false negative problem, it employs a symmetric, denoised contrastive loss based on NCE, which uses the known interaction graph to deterministically remove in-batch false negatives, reducing intra-class repulsion. Second, guided by our ”magnitude-as-noise” hypothesis, it projects all embeddings onto a unit hypersphere via L2 normalization. This structurally removes the popularity signal to isolate the preference encoded in the vector’s direction. Finally, we demonstrate that a model built on these principles naturally resolves signal ambiguity, learning to anchor its alignment not on superficial sentiment but on objective, information-rich features.
Our main contributions are:
@@
  • A Unified Geometric Framework (SymCERE): We propose a method that structurally mitigates false negatives, popularity bias, and signal ambiguity. It achieves this by symmetrically applying a denoised NCE loss for cross-modal alignment and integrating geometric debiasing via L2 normalization.
@@
  • Extensive Empirical Validation: We demonstrate the effectiveness of SymCERE through comprehensive experiments on 15 diverse datasets, achieving significant improvements (e.g., up to 43.6% in NDCG@10) over strong baselines.
@@
  • Discovery of Semantic Anchoring: We find, through in-depth LIME analysis, that our model learns to prioritize product-specific terminology (e.g., ‘OEM’, ‘compatible’) over generic sentiment (e.g., ‘good’, ‘nice’). We term this mechanism ‘Semantic Anchoring’, providing evidence that robust multi-modal alignment is driven by factual product attributes.
@@ 2. Related Work
This work lies at the intersection of multi-modal recommendation, Graph Neural Networks (GNNs), and contrastive learning (CL). GNNs like LightGCN (He et al., 2020) form the backbone of many modern recommenders by modeling the user-item interaction graph. To enhance the learned representations, contrastive learning (CL) has been widely adopted, typically using the InfoNCE loss with augmented data (Wu et al., 2021). However, this common approach may overlook the potential of CL to address deeper, interconnected geometric problems.
@@ 2.1. A Triad of Interconnected Challenges
We frame the primary obstacles in review-enhanced recommendation not as isolated issues requiring bespoke modules, but as an interconnected triad that results in a ”Fusion Gap.”
2.1.1. The False Negative Problem
The standard InfoNCE loss has a key limitation: by treating all non-positive samples as negatives, it inherently penalizes ”false negatives”—semantically similar items that are not the anchor’s direct positive pair (Huynh et al., 2022; Chen et al., 2022). This ”intra-class repulsion” negatively affects the embedding space by pushing similar items apart. While heuristic correction methods exist (Chen et al., 2022), they can be complex. In recommendation, we have the advantage of a known interaction history, which allows for more deterministic mitigation strategies.
2.1.2. Popularity Bias as a Geometric Artifact
Popularity bias, the over-recommendation of popular items, is a known challenge to fairness and diversity (Klimashevskaia et al., 2024). We approach this from a geometric perspective, positing that popularity is often encoded as the magnitude (L2-norm) of an item’s embedding vector (Lin et al., 2022). This magnitude acts as a noisy signal that obscures user preferences. Therefore, we treat L2 normalization not just as a technical step, but as a debiasing mechanism that structurally removes the popularity signal to encourage the model to encode preference solely in the vector’s direction (Wang and Isola, 2020).
2.1.3. Signal Ambiguity in Multi-Modal Fusion
Fusing collaborative signals with review texts enriches representations but introduces ”signal ambiguity”. A review conflates objective product attributes (e.g., ”OEM part,” ”compatible gasket”) with subjective sentiment (e.g., ”good,” ”great”) (Wu et al., 2014). A simple alignment may lead to learning spurious correlations from generic sentiment. We hypothesize that a robustly formulated CL method can resolve this ambiguity by forcing alignment based on stable, objective features, effectively using reviews as ”soft metadata.”
@@ 2.2. Our Approach: A Unified Solution via Principled CL
The literature often presents a range of methods that address these issues individually for denoising, debiasing, and multi-modal fusion (Lou et al., 2024; Lin et al., 2023, 2024). We diverge from this trend, arguing that modifying the core CL objective itself offers a more direct and unified solution. To this end, we build upon NCE, a supervised contrastive loss that addresses the limitations of the InfoNCE objective by reformulating it to eliminate intra-class repulsion. By integrating a symmetric NCE loss with geometric debiasing (L2 normalization), SymCERE is designed to simultaneously mitigate false negatives and popularity bias, thereby resolving signal ambiguity.
@@ 3. Methodology
Our proposed framework, SymCERE, aligns collaborative signals (graph modality) and semantic information (text modality) into a unified hyperspherical space. The overall architecture is illustrated in Figure 2.
Figure 2. The SymCERE Framework. The model bridges the Fusion Gap through two key geometric mechanisms: (1) L2 Normalization projects embeddings onto a hypersphere to structurally neutralize popularity bias (magnitude noise). (2) A Symmetric NCE Loss utilizes the known interaction mask $M$ to deterministically exclude false negatives (grey cells) from the denominator, preventing intra-class repulsion. This unified objective aligns collaborative signals ($\mathbf{g}$) and semantic signals ($\mathbf{t}$) while preserving ranking order via BPR.
@@ 3.1. Modality Encoders
3.1.1. Graph Encoder (Collaborative Signals).
We employ established GNN architectures as the graph encoder, $f_{\text{graph}}(\cdot)$. In this study, we utilize both LightGCN (He et al., 2020) (a linear model) and NGCF (Wang et al., 2019) (a non-linear model) as backbones to evaluate the framework’s versatility.
LightGCN refines embeddings $\mathbf{E}^{(0)}$ over $K$ layers via linear propagation. The final representation is a weighted combination of all layers:
(1) $\mathbf{g}_{u}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{u}^{(k)},\quad\mathbf{g}_{i}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{i}^{(k)}.$
where $\alpha_{k}=1/(K+1)$. Regardless of the backbone, the interaction $(u,v)$ representation is computed as $\mathbf{g}_{u,v}=(\mathbf{g}_{u}+\mathbf{g}_{v})/2$. This interaction-centric representation serves as the target for cross-modal alignment.
3.1.2. Text Encoder (Semantic Signals).
A pre-trained LLM adapted with LoRA serves as the text encoder, $f_{\text{text}}(\cdot)$. Given review text $r_{u,v}$ for interaction $(u,v)$, the encoder produces a dense semantic embedding $\mathbf{t}_{u,v}=f_{\text{text}}(r_{u,v})$. To ensure temporal integrity and prevent leakage, $r_{u,v}$ consists strictly of the review text written by user $u$ for item $v$ at the time of interaction. We treat this text as a source of static, intrinsic product attributes (”Semantic Anchoring”) rather than dynamic sentiment, facilitating robust alignment.
3.2. Unified Contrastive Alignment Framework
SymCERE addresses the geometric distortions caused by popularity bias and contrastive false negatives through a principled fusion objective.
3.2.1. Prerequisite: L2 Normalization for Stability.
To mitigate popularity bias, which often manifests as magnitude artifacts (Lin et al., 2022), we enforce L2 normalization on all embeddings before loss calculation. This projects representations onto a unit hypersphere, ensuring alignment is based purely on directional similarity:
(2) $\hat{\mathbf{g}}_{u,v}=\frac{\mathbf{g}_{u,v}}{\|\mathbf{g}_{u,v}\|_{2}},\quad\hat{\mathbf{t}}_{u,v}=\frac{\mathbf{t}_{u,v}}{\|\mathbf{t}_{u,v}\|_{2}}.$
3.2.2. Multi-Task Contrastive Objectives.
We combine cross-modal alignment with intra-modal regularization.
@@ 3.2. Unified Contrastive Alignment Framework
Cross-Modal Alignment with Symmetric NCE Loss ($\mathcal{L}_{\text{cross-modal}}$).
Standard InfoNCE loss suffers from the false negative problem, where unobserved but valid positive items are treated as negatives, causing intra-class repulsion. To resolve this, we propose a Symmetric NCE Loss based on Noise-Contrastive Estimation (Gutmann and Hyvärinen, 2010). This objective modifies the standard contrastive formulation to deterministically exclude false negatives.
In recommendation, the full interaction history $\mathcal{R}$ is known. For a mini-batch $\mathcal{B}$, we construct a ”true negative” set $\mathcal{N}_{i}$ for an anchor interaction $i=(u_{i},v_{i})$ by excluding any sample $j$ where the user $u_{i}$ has a known interaction with item $v_{j}$:
(3) $\mathcal{N}_{i}=\{j\in\mathcal{B}\mid j\neq i\land(u_{i},v_{j})\notin\mathcal{R}\}.$
This ensures that the denominator only penalizes items that are truly unobserved for the user. The Graph-to-Text alignment loss ($\mathcal{L}_{G\to T}$) is defined as:
(4) $\begin{split}\mathcal{L}_{G\to T}=-\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\Big[&\frac{\hat{\mathbf{g}}_{u_{i},v_{i}}^{\top}\hat{\mathbf{t}}_{u_{i},v_{i}}}{\tau}\\ &-\log\Big(\sum_{k\in\{i\}\cup\mathcal{N}_{i}}\exp\frac{\hat{\mathbf{g}}_{u_{i},v_{i}}^{\top}\hat{\mathbf{t}}_{u_{k},v_{k}}}{\tau}\Big)\Big].\end{split}$
Symmetrically, the Text-to-Graph loss ($\mathcal{L}_{T\to G}$) is computed by swapping the roles of anchor and positive, ensuring bidirectional alignment. The final cross-modal loss is:
(5) $\mathcal{L}_{\text{cross-modal}}=\frac{1}{2}\left(\mathcal{L}_{G\to T}+\mathcal{L}_{T\to G}\right).$
Intra-Modal Enhancement ($\mathcal{L}_{\text{intra-modal}}$).
To improve the robustness of individual modalities, we employ standard InfoNCE loss on augmented views (edge dropout for graph $\hat{\mathbf{g}}^{\prime}$, word masking for text $\hat{\mathbf{t}}^{\prime}$):
(6) $\mathcal{L}_{\text{intra-modal}}=\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{g}},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{InfoNCE}}(\hat{\mathbf{t}},\hat{\mathbf{t}}^{\prime}).$
@@
We further regularize the graph embedding space using the Bayesian Personalized Ranking (BPR) loss to enforce relative preference ordering:
(7) $\mathcal{L}_{\text{BPR}}=\sum_{(u,v^{+},v^{-})\in\mathcal{D}}-\log\sigma(\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{v^{+}}-\hat{\mathbf{g}}_{u}^{\top}\hat{\mathbf{g}}_{v^{-}}).$